OpenStudy (anonymous):

x=w+y+z/5 solve for y

OpenStudy (anonymous):

y=-z/5 +x-w

OpenStudy (anonymous):

how did you get that answer. can you show how u set it up please?

OpenStudy (anonymous):

ok wait one moment

OpenStudy (anonymous):

y +w+z/5 = x subtract (z/5) +w on both sides of the expression y=-z/5 +x-w

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

lol ok